OpenStudy (rayman5766):

Simplify. –20 + (–2) • (–6) =

OpenStudy (rayman5766):

Simplify. –20 + (–2) • (–6) = A. –132 B. –32 C. –8 D. 132

OpenStudy (akashdeepdeb):

Multiplication rules: -a * b = -ab -a * -b = ab a * b = ab a * -b = -ab Additions rules: -a + (-b) = -a-b -a + (b) = = -a+ b = b - a a + (-b) = a-b a + (b) = a + b NOTE: Subtraction is just addition of the negative element.

OpenStudy (rayman5766):

oh ok I get it now